the odds in favor of Trudy beating her friend in a round of golf are 1:9. Find the probability that Trudy will lose.
Answer by stanbon(75887)

the odds in favor of Trudy beating her friend in a round of golf are 1:9. Find the probability that Trudy will lose.
-----------
Formula: "odds FOR event A" = P(A)/P(not A)
--------
Your Problem:
1:9 = P(A)/P(not A)
P(A) = (1/9)*P(not A)
----------
But P(A) + P(not A) = 1
So, P(not A) = 1 - P(A)
And P(not A) = 1 - [(1/9)P(not A)]
(10/9)P(not A) = 1
P(not A) = 9/10
=======================
That was a bit long-winded. Here's a shorter version.
Since odds FOR = 1:9
P(A) = 1/(1+9) = 1/10
So P(A') = 9/10
====================
Cheers,
Stan H.